# Automatically Identifying Local and Global Circuits with Linear Computation Graphs

### v1

@@ Abstract
Circuit analysis of any certain model behavior is a central task in mechanistic interpretability. We introduce our circuit discovery pipeline with sparse autoencoders (SAEs) and a variant called skip SAEs. With these two modules inserted into the model, the model’s computation graph with respect to OV and MLP circuits becomes strictly linear. Our methods do not require linear approximation to compute the causal effect of each node. This fine-grained graph enables identifying both end-to-end and local circuits accounting for either logits or intermediate features. We can scalably apply this pipeline with a technique called Hierarchical Attribution. We analyze three kind of circuits in GPT2-Small, namely bracket, induction and Indirect Object Identification circuits. Our results reveal new findings underlying existing discoveries.
@@ 1 Introduction
Recent years have seen the rapid progress of mechanistically reverse engineering Transformer language models (Vaswani et al., 2017). Conventionally, researchers seek to find out how neural networks organize information in its hidden activation space (Olah et al., 2020a; Gurnee et al., 2023; Zou et al., 2023) (i.e. features) and how learnable weight matrices connect and (de)activate them (Olsson et al., 2022; Wang et al., 2023; Conmy et al., 2023) (i.e. circuits). One fundamental problem of studying attention heads and MLP neurons as interpretability primitives is their polysemanticity, which under the assumption of linear representation hypothesis is mostly due to superposition (Elhage et al., 2022; Larson, 2023; LaurenGreenspan & keith_wynroe, 2023). Thus there is no guarantee to explain how these components impact model behavior out of the interested distribution. Additionally, circuit analysis based on attention heads is coarse-grained in that it lacks effective methods to explain the intermediate activations.
@@
Sparse Autoencoders (SAEs) (Bricken et al., 2023; Cunningham et al., 2023) have shown its potential in extracting features from superposition in an unsupervised manner. This opens up a new perspective of understanding model internals by interpreting the activation of SAE features. It also poses a natural research question: how to gracefully leverage SAEs for circuit analysis? Compared to prior work along this line (Cunningham et al., 2023; He et al., 2024; Marks et al., 2024), our main contributions are as follows.
Figure 1: Overview of our method. For a given input, we (1) run forward process once with MLP computation replaced by Skip SAEs. (2) Then a subgraph is isolated for a given input with Hierarchical Attribution in one backward. (3) We then interpret important QK attention involved in the identified circuit.
  • We propose to utilize Skip SAEs, a variant of Sparse Autoencoders to sparsely approximate the computation of MLP layers. This extends the linear analysis of Transformer circuits (Elhage et al., 2021; He et al., 2024).
  • For a given input, OV + Skip SAE (i.e. MLP) circuits strictly form a Linear Computation Graph without linear approximation of any non-linear function. This precious linearity enables circuit discovery and evaluation with only one forward and one backward.
@@
  • We present a specific example that our analysis offers more detailed insight into how each single SAE feature contributes to a desired behavior e.g. forms a crucial QK attention, or linearly activate a subsequent node in the computation graph. Such observations are not reported by existing work studying circuits in coarser granularity.
@@ 2.1 Sparse Autoencoder Features as Analytic Primitives
Sparse Autoencoder (SAE) is a recently emerging method to take features of model activation out of superposition (Elhage et al., 2022). Existing work has suggested empirical success in the interpretability of SAE features with respect to both human evaluation (Bricken et al., 2023) and automatic evaluation (Bills et al., 2023).
@@
We train Sparse Autoencoders to decompose all modules that write into the residual stream (i.e. Word Embedding, Attention output and MLP output). Then we can derive how a residual stream activation is composed of SAE features:
$x=\sum_{S\in\text{Upstream SAEs}}\left(\sum_{i=1}^{d_{\text{SAE}}}f_{i}^{S}{W_{D}^{S}}_{i}+\varepsilon^{S}\right)+p,$ (2)
@@
where $f_{i}^{S}$ and $\varepsilon^{S}$ are feature activation and SAE error term of each upstream SAEs. $p$ is the positional embedding of current token. Since all submodules read and write into the residual stream, such partition is crucial to connect upstream SAE features to downstream ones.
2.2 Tackling MLP Non-linearity with Skip SAEs
(a) Skip SAEs operate both computation and sparse coding.
(b) Skip SAEs improve reconstruction and sparsity in 8 out of 12 layers.
Figure 2: Skip SAEs take in the same input as MLP layers and are trained to reconstruct MLP outputs. For any given input, Skip SAEs serve as a surrogate of MLP and simultaneously map the output to a linear sparse coding. In comparison, linear analysis is intractable or requires approximation for simply reconstructing the original MLP output (Figure 2(a)). Reconstruction performance and sparsity of Skip SAEs is no worse than original MLP SAEs (Figure 2(b)).
@@ 2.2 Tackling MLP Non-linearity with Skip SAEs
The denseness and non-linearity of MLP in Transformers makes it difficult for sparse attribution of MLP features. The key difficulty lies in that MLP activation functions have a privileged basis (Elhage et al., 2023). Computation of MLP non-linearity must go through such an orthogonal basis of the MLP hidden space. There is no guarantee to observe sparse and informative correspondence between MLP neurons and learned SAE features. This annoying non-linearity cuts off connection of earlier-layer SAE features and MLP output (with respect to linear algebraic operations).
@@
To tackle this problem, we develop a new method called Skip SAEs to get round the MLP non-linearity. Skip SAEs are generalized form of SAEs, which decouple input and output of SAEs and allow predicting future activations given an earlier activation of the model. As shown in Figure 2(a), Skip SAEs take in the pre-MLP activation and yield a sparse decomposition of MLP output.
Why are Skip SAEs linear even though they have non-linearity?
One useful thought experiment is to ignore the de-activated SAE features, with respect to a given input, as if they do not exist. This does not change the computation of SAEs but will turn it into a sparse and linear operation. Though this trick can also be applied to MLPs, the key difference in interpretability is that SAEs have an interpretable basis. Strictly speaking, what we actually claim is that the sparse decomposition of MLP output is a linear function of MLP input. This is also compatible with conventional SAEs.
Intuitively, Skip SAEs operate both computation and sparse coding. This raises a natural practical concern of Skip SAEs that they might introduce more SAE error and thus harm interpretability. We conduct a series of experiments to test this hypothesis. Results in Figure 2(b) indicate that surprisingly, Skip SAEs achieve better sparsity and reconstruction than their counterparts trained to merely recover MLP outputs.
@@ 2.3 QK and OV Circuits Are Independent Linear Operators on SAE Features
QK and OV circuits account for how tokens attend to one another, and how information passes to downstream layer, respectively. The linearity and independence of these two components have been widely discussed in previous work (Elhage et al., 2021; He et al., 2024). Specifically, QK circuits serve as a bilinear operator of any two residual streams i.e. $x_{i}{W_{Q}^{h}}^{T}W_{K}^{h}x_{j}^{T}$ for token $i$ and $j$. OV circuits operates a linear transformation to all past residual streams and take a weighted sum of them.
What about Softmax?
We follow Elhage et al. (2021) to first fix Softmax output (i.e. attention pattern) as constant. When analysis on OV + Skip SAE circuit is finished, we come back to identify the core contributors to any specific pre-softmax entry. This two-step process does not interpret the attention head as a whole but instead treats QK and OV as two sub-problems.
@@ 3 Isolating Interpretable Circuits with Hierarchical Attribution
Figure 3: Our Hierarchical Attribution detaches unrelated nodes immediately after they receive gradient and stops their back propagation. While standard attribution detaches nodes after the backward process is completed. (Figure 3(a)). We sweep the number of remaining nodes i.e. sparsity and compare the logit recovery i.e. faithfulness of the identified subgraph. Experiments are conducted on 20 IOI samples (See Section 5) across 30 sparsity thresholds. Results in Figure 3(b) show that Hierarchical Attribution consistently outperforms standard attribution.
Now we have obtained a linear computation graph including all OV and MLP modules, reflecting the model’s internal information flow. We introduce in this section how to isolate and evaluate a subgraph of the key SAE features related to any interested output.
@@
We are given a linear computation graph $G=(V,E)$, which is a directed acyclic graph with only one root node (i.e. logit). Each node $v\in V$ refers to an activated feature in model forward process. The node weight $a_{v}$ refers to the activation of node $v$. Each edge $v\to u\in E$ represents that $a_{v}$ linearly affects $a_{u}$ by the edge weight $k_{v,u}$. For any non-leaf node $u$, its activation is completely determined by its direct predecessors i.e. $a_{u}=\sum_{v\to u\in E}k_{v,u}a_{v}$
@@
We denote word embedding SAE features and the position embedding as interpretable leaf nodes [^1] We notice that not all SAE features are interpretable. We adopt a series of methods to furthur improve the interpretability of SAEs. See Appendix A. SAE errors also have zero in-degree but we cannot establish any explanation on these nodes. Thus we call them uninterpretable leaf nodes.
@@
To identify a subgraph accounting for the desired output, we prune a number of unrelated nodes in the original linear computation graph.
@@
The operation of detaching a node $v$ from graph $G$ is to get a induced subgraph $G^{\prime}=G[V/v]$, which removes $v$ and all edges connecting to $v$ from $G$.
We first need to detach all SAE errors since they cannot be interpreted, despite their empirically positive correlation to model performance (wesg, 2024). In the rest of the graph with all leaf nodes are interpretable leaf nodes, we need to detach nodes unrelated to the task.
@@
For graphs with a small number of nodes, a simple solution is to manually inspect the interpretation of SAE features and their causal relation. This is often useful in understanding local behaviors but may be labor-extensive at scale.
@@
A natural idea would be running backward once and detaching nodes with $\operatorname{attr}_{v,t}$ lower than a given threshold $\tau$, as adopted in most prior work (Conmy et al., 2023; Marks et al., 2024). We propose to operate a breadth first search style attribution pipeline we call Hierarchical Attribution.
Hierarchical Attribution detaches nodes on backward process instead of after backward, as shown in Figure 3(a) and a pseudo-code implementation in Appendix C. When performing model backward, we stop the gradient propagation of any node $v$ that has $\operatorname{attr}_{v,t}<\tau$. This affects the attribution score of all predecessors of $v$. After we finish the backward propagation, all nodes with gradient make up our desired subgraph. Intuitively, attribution through detached nodes should not be taken into account otherwise their effect depends on excluded node in the final subgraph.
@@
This theorem allows us to obtain how much $G^{\prime}=G[V/v]$ accounts for the root node activation instantly after we finish the pruning. Another advantage of such evaluation besides efficiency lies in that it derives the causal effect of circuits without any intervention in the forward process. It saves circuit evaluation from backup behaviors (Wang et al., 2023) (also known as hydra effects (McGrath et al., 2023)) due to ablation.
In Figure 3(b), we empirically validate the advantage of Hierarchical Attribution over standard attribution method in Indirect Object Identification circuit discovery (Wang et al., 2023).

### v2

@@ Abstract
Circuit analysis of any certain model behavior is a central task in mechanistic interpretability. We introduce our circuit discovery pipeline with Sparse Autoencoders (SAEs) and a variant called Transcoders. With these two modules inserted into the model, the model’s computation graph with respect to OV and MLP circuits becomes strictly linear. Our methods do not require linear approximation to compute the causal effect of each node. This fine-grained graph identifies both end-to-end and local circuits accounting for either logits or intermediate features. We can scalably apply this pipeline with a technique called Hierarchical Attribution. We analyze three kinds of circuits in GPT-2 Small: bracket, induction, and Indirect Object Identification circuits. Our results reveal new findings underlying existing discoveries.
@@ 1 Introduction
Recent years have seen the rapid progress of mechanistically reverse engineering Transformer language models (Vaswani et al., 2017). Conventionally, researchers seek to find out how neural networks organize information in its hidden activation space (Olah et al., 2020a; Gurnee et al., 2023; Zou et al., 2023) (i.e. features) and how learnable weight matrices connect and (de)activate them (Olsson et al., 2022; Wang et al., 2023; Conmy et al., 2023) (i.e. circuits). One fundamental problem of studying attention heads and MLP neurons as interpretability primitives is their polysemanticity, which under the assumption of linear representation hypothesis is mostly due to superposition (Elhage et al., 2022; Larson, 2023; LaurenGreenspan & keith_wynroe, 2023). Thus, there is no guarantee of explaining how these components impact model behavior out of the interested distribution. Additionally, circuit analysis based on attention heads is coarse-grained because it lacks effective methods to explain the intermediate activations.
@@
Sparse Autoencoders (SAEs) (Bricken et al., 2023; Cunningham et al., 2023) have shown their potential in extracting features from superposition in an unsupervised manner. This opens up a new perspective of understanding model internals by interpreting the activation of SAE features. It also poses a natural research question: how to gracefully leverage SAEs for circuit analysis? Compared to prior work along this line (Cunningham et al., 2023; He et al., 2024; Marks et al., 2024), our main contributions are as follows.
Figure 1: Overview of our method. For a given input, we (1) run forward pass once with MLP computation replaced by Trans. (2) Then a subgraph is isolated for a given input with Hierarchical Attribution in one backward. (3) We then interpret important QK attention involved in the identified circuit.
  • We propose to utilize Transcoders, a variant of Sparse Autoencoders, to sparsely approximate the computation of MLP layers. This extends the linear analysis of Transformer circuits (Elhage et al., 2021; He et al., 2024).
  • For a given input, OV + Transcoder (i.e., MLP) circuits strictly form a Linear Computation Graph without linear approximation of any non-linear function. This precious linearity enables circuit discovery and evaluation with only one forward and one backward.
@@
  • We present a specific example in our analysis that offers more detailed insight into how each single SAE feature contributes to a desired behavior, e.g., forms a crucial QK attention or linearly activates a subsequent node in the computation graph. Such observations are not reported by existing work studying circuits in coarser granularity.
@@ 2.1 Sparse Autoencoder Features as Analytic Primitives
Sparse Autoencoder (SAE) is a recently emerging method to take features of model activation out of superposition (Elhage et al., 2022). Existing work has suggested empirical success in the interpretability of SAE features concerning both human evaluation (Bricken et al., 2023) and automatic evaluation (Bills et al., 2023).
@@
We train Sparse Autoencoders on GPT-2 (Radford et al., 2019) to decompose all modules that write into the residual stream (i.e. Word Embedding, Attention output and MLP output). Then, we can derive how a residual stream activation is composed of SAE features:
$x=\sum_{\mathcal{S}\in\text{Upstream SAEs}}\left(\sum_{i=1}^{d_{\text{SAE}}}f_{i}^{\mathcal{S}}{W_{D}^{\mathcal{S}}}_{i}+\varepsilon^{\mathcal{S}}\right)+p,$ (2)
@@
where $f_{i}^{\mathcal{S}}$ and $\varepsilon^{\mathcal{S}}$ are feature activation and SAE error term of each upstream SAE $\mathcal{S}$. $p$ is the positional embedding of the current token. Since all submodules read and write into the residual stream, such a partition is crucial to connect upstream SAE features to downstream ones.
2.2 Tackling MLP Non-linearity with Transcoders
@@
The denseness and non-linearity of MLP in Transformers make sparse attribution of MLP features difficult. Since MLP activation functions have a privileged basis (Elhage et al., 2023), computation of MLP non-linearity must go through such an orthogonal basis of the MLP hidden space. There is no guarantee of observing sparse and informative correspondence between MLP neurons and learned SAE features. This annoying non-linearity cuts off the connection of upstream SAE features and MLP output (with linear algebraic operations).
@@
To tackle this problem, we develop a new method called Transcoders to get around the MLP non-linearity. Transcoders are generalized forms of SAEs, which decouple the input and output of SAEs and allow for predicting future activations given an earlier model activation. Transcoders take in the pre-MLP activation and yield a sparse decomposition of MLP output. Formally, a Transcoder and its optimization objective can be written as:
$f$ $=\operatorname{ReLU}(W_{E}x+b_{E})$ (3)
$\hat{y}$ $=W_{D}f$
$\mathcal{L}$ $=\lVert y-\hat{y}\rVert_{2}^{2}+\lambda\lVert f\rVert_{1},$
which only differs from those of an SAE (Eq. 1) by the label activation $y\in\mathbb{R}^{d_{\text{model}}}$ unbound with input activation $x$.
Key difference between Transcoders and MLP
We may find Transcoders and MLP with similar architecture: both are two fully connected blocks interspersed with an activation function. It’s natural to ask why the non-linear activation function in MLP is deemed as an obstacle in circuit analysis but that in Transcoders is allowed. The key difference is that by constraining the sparsity, Transcoders neurons (which are just features) have an interpretable basis. When computing how upstream feature $f_{i}^{\mathcal{S}}$ contributes to activated downstream feature $f_{j}^{\mathcal{T}}$ of Transcoder $\mathcal{T}$, it holds that $f_{j}^{\mathcal{T}}=f_{i}^{\mathcal{S}}\left(W_{E}^{\mathcal{T}}W_{D}^{\mathcal{S}}\right)_{ji}$. The $\left(W_{E}^{\mathcal{T}}W_{D}^{\mathcal{S}}\right)_{ji}$ part remains constant across inputs, which leads to an edge invariance between upstream and downstream features.
Intuitively, this means when a main upstream contributor to a downstream feature has been activated in a different input, we can largely expect this downstream feature to be activated again unless some new resistances (upstream features with negative edges) have also been introduced.
In contrast, we cannot find such invariant edges through MLP. Any connection from upstream to MLP output is indefinite, so we could only find linear approximations to measure these connections under local changes.
@@ 2.3 QK and OV Circuits Are Independent Linear Operators on SAE Features
QK and OV circuits account for how tokens attend to one another and how information passes to downstream layers, respectively. The linearity and independence of these two components have been widely discussed in previous work (Elhage et al., 2021; He et al., 2024). Specifically, QK circuits serve as a bilinear operator of any two residual streams w.r.t token $i$ and $j$:
$\operatorname{AttnScore}^{h}(x)_{ij}$ $=x_{i}{W_{Q}^{h}}^{T}W_{K}^{h}x_{j}^{T}$ (4)
$=\sum_{\mathcal{S},\mathcal{T}\in\text{Upstream SAEs}}\sum_{p=1}^{d_{\text{SAE}}}\sum_{q=1}^{d_{\text{SAE}}}f_{i,p}^{\mathcal{S}}{W_{D}^{\mathcal{S}}}_{p}{W_{Q}^{h}}^{T}W_{K}^{h}{W_{D}^{\mathcal{T}}}_{q}^{T}f_{j,q}^{\mathcal{T}},$
where $f_{i,p}$ means the activation of the feature $p$ at token $i$, and $W_{Q}^{h},W_{K}^{h}$ are a given head $h$’s the query and key transformation. This decomposition shows how every pair of upstream features contributes to the attention score, making tokens containing critical information get attended.
Once the attention score is determined, we can then move on to the OV circuits, which apply a linear transformation to all past residual streams and take a weighted sum:
$\operatorname{Attn}(x)_{i}$ $=\sum_{h}\operatorname{AttnOutput}^{h}(x)_{i}$ (5)
$=\sum_{h}\sum_{j}{\color[rgb]{0.25,0.41,0.88}\operatorname{AttnPattern}^{h}(x)_{i,j}}W_{O}^{h}W_{V}^{h}x_{j},$
where $W_{O}^{h},W_{V}^{h}$ are a given head $h$’s output and value transformation. With AttnPattern determined in the QK circuits, how upstream features affect downstream are successively determined since $W_{O}^{h}W_{V}^{h}$ is invariant.
From an input-independent perspective, the quadratic coefficient ${W_{D}^{\mathcal{S}}}_{p}{W_{Q}^{h}}^{T}W_{K}^{h}{W_{D}^{\mathcal{T}}}_{q}$ shows how feature pairs co-work for every attention score. Then, ${W_{E}^{\mathcal{S}}}_{p}W_{O}^{h}W_{V}^{h}{W_{D}^{\mathcal{T}}}_{q}$ (obtained by adding SAE encoder and decoder terms to Eq. 5) determines the edge connecting upstream features and attention output features under a specific attention pattern. This two-step paradigm gives us a simplified and feature-based version of attention functionality and allows a fine-grained analysis through attention in a non-approximated manner.
In real-world applications, we often want to attribute an interested output (e.g., logits) to filter out critical features, which is a backward procedure. For the sake of a linear and exact attribution result, we can reverse the above two-step paradigm and 1) attribute through OV + Transcoder circuits and then 2) select important attention, attribute its attention score through the current QK and once again the upstream OV + Transcoder circuits (showed in Figure. 2(a)). The second step may be repeated several times to attribute attention important to another attention.
@@ 3 Isolating Interpretable Circuits with Hierarchical Attribution
Figure 2: Our Hierarchical Attribution detaches unrelated nodes immediately after they receive gradient and stops their backpropagation, while standard attribution detaches nodes after the backward pass is completed. (Figure 2(a)). We sweep the number of remaining nodes, i.e., sparsity, and compare the logit recovery, i.e., faithfulness of the identified subgraph. Experiments are conducted on 20 IOI samples (See Section 5) across 30 sparsity thresholds. Results in Figure 2(b) show that Hierarchical Attribution consistently outperforms standard attribution.
We have now obtained a linear computation graph including all OV and MLP modules, reflecting the model’s internal information flow. This section introduces how to isolate and evaluate a subgraph of the key SAE features related to any interested output.
@@
We are given a linear computation graph $G=(V,E)$, which is a directed acyclic graph. Each node $v\in V$ refers to an activated feature in the model forward pass. The node weight $a_{v}$ refers to the activation of node $v$. Each edge $v\to u\in E$ represents that $a_{v}$ linearly affects $a_{u}$ by the edge weight $k_{v,u}$. For any non-leaf node $u$, its activation is completely determined by its direct predecessors, i.e., $a_{u}=\operatorname{ReLU}\left(\sum_{v\to u\in E}k_{v,u}a_{v}\right)$.
The term linear computation graph means every edge in the graph represents a linear function (under fixed attention scores). This guarantees a one-hop linear effect of activated features. It’s not necessary that indirect effects between any two nodes are still linear since we allow a ReLU gate inside the nodes, stopping unactivated nodes from forwarding further.
@@
We denote word embedding SAE features and the position embedding as interpretable leaf nodes [^1] We notice that not all SAE features are interpretable. We adopt a series of methods to improve the interpretability of SAEs further. See Appendix A. SAE errors also have zero in degree, but we cannot establish any explanation for these nodes. Thus, we call them uninterpretable leaf nodes.
@@
We prune unrelated nodes in the original linear computation graph to identify a subgraph accounting for the desired output.
@@
The operation of detaching a node $v$ from graph $G$ is to get an induced subgraph $G^{\prime}=G[V/v]$, which removes $v$ and all edges connecting to $v$ from $G$.
We first need to detach all SAE errors since they cannot be interpreted, despite their empirically positive correlation to model performance (Gurnee, 2024). In the rest of the graph, with all leaf nodes being interpretable leaf nodes, we need to detach nodes unrelated to the task.
@@
For graphs with a small number of nodes, a simple solution is to manually inspect the interpretation of SAE features and their causal relation. This is often useful in understanding local behaviors but may be labor-intensive at scale.
@@
A natural idea would be running backward once and detaching nodes with $\operatorname{attr}_{v,t}$ lower than a given threshold $\tau$, as adopted in most prior work (Conmy et al., 2023; Marks et al., 2024). We propose to operate a breadth-first search style attribution pipeline we call Hierarchical Attribution.
Hierarchical Attribution detaches nodes on backward pass instead of after backward, as shown in Figure 2(a) and a pseudo-code implementation in Appendix C. When performing model backward, we stop the gradient propagation of any node $v$ that has $\operatorname{attr}_{v,t}<\tau$. This affects the attribution score of all predecessors of $v$. After we finish the backward propagation, all nodes with gradients make up our desired subgraph. Intuitively, attribution through detached nodes should not be taken into account; otherwise, their effect depends on excluded nodes in the final subgraph.
@@
This theorem allows us to instantly obtain how much $G^{\prime}=G[V/v]$ accounts for the root node activation after we finish the pruning. Besides efficiency, another advantage of such evaluation is that it derives the causal effect of circuits without any intervention in the forward pass. It saves circuit evaluation from backup behaviors (Wang et al., 2023) (also known as hydra effects (McGrath et al., 2023)) due to ablation.
In Figure 2(b), we empirically validate the advantage of Hierarchical Attribution over the standard attribution method in Indirect Object Identification circuit discovery (Wang et al., 2023).